# Tractable Weighted First-Order Model Counting with Bounded Treewidth Binary Evidence

### v1

@@ 1 Introduction
Evidence is the interpretation of certain predicates. For example, consider the sentence
@@
over the domain $\Delta=\{1,2,3,4\}$. An evidence can be the specification that $E(1,2),E(1,3),E(2,3),E(1,4)$ and their reverse atoms are true and other atoms of $E$ are false. In this case, each model is an assignment of $I(1),\cdots,I(4)$ satisfying $\Psi$, which represents an independent set of the graph with vertices $\{1,2,3,4\}$ and undirected edges $(1,2),(1,3),(2,3),(1,4)$. The WFOMC of $\Psi$ with the above evidence over $\Delta$ with unit weight for each model equals to the number of independent sets on the graph.
@@
In this work, we study the binary evidence whose Gaifman graph, the undirected graph with vertices corresponding to domain elements in which two vertices $a$ and $b$ are connected if and only if there is a binary predicate $P$ such that the truth value of $P(a,b)$ or $P(b,a)$ is specified in the binary evidence, is of bounded treewidth. Our primary result is to show that computing WFOMC for $\text{FO}^{2}$ and $\text{C}^{2}$ with such binary evidence can be done in time polynomial in the domain size. Our approach applies to counting problems in combinatorics on bounded-treewidth graphs by encoding the problem by WFOMC of a fixed $\text{FO}^{2}$ or $\text{C}^{2}$ sentence with such binary evidence.
@@ 1.1 Our Contributions
We propose an algorithm to compute WFOMC for an $\text{FO}^{2}$ or $\text{C}^{2}$ sentence with cardinality constraints, unary evidence and binary evidence whose Gaifman graph is of bounded treewidth. The algorithm runs in time polynomial in the domain size. This offers a new approach to overcome the symmetry limitation in lifted inference, and expands the expressive power and applicability of WFOMC. The key idea of our algorithm is the dynamic programming for computing WFOMC for the universally quantified $\text{FO}^{2}$ sentence with evidence on the tree decomposition of the underlying Gaifman graph. Existing techniques help to extend the result to $\text{FO}^{2}$ and $\text{C}^{2}$ with cardinality constraints.
@@ 2.5 Evidence
The two forms of the evidence can be transformed to each other. Transforming the closed-world representation to the open-world representation is straightforward by adding the default negative ground literals to the evidence set. For instance, the closed-world binary evidence $\mathcal{E}=\{R(1,2),R(2,2)\}$ of the predicate set $\mathcal{P^{\prime}}=\{R\}$ over the domain $\Delta=\{1,2\}$ corresponds to the open-world binary evidence $\mathcal{E}^{\prime}=\{\lnot R(1,1),R(1,2),\lnot R(2,1),R(2,2)\}$. On the other hand, for example, consider the open-world binary evidence $\mathcal{E}=\{R(1,2),R(2,3),\lnot R(3,4)\}$. To transform $\mathcal{E}$ to the closed-world form, we introduce two fresh binary predicates $R^{\top}$ and $R^{\bot}$, add the sentence $\forall x\forall y:(R^{\top}(x,y)\to R(x,y))\land(R^{\bot}(x,y)\to\lnot R(x,y))$ and write $\mathcal{E}^{\prime}=\{R^{\top}(1,2),R^{\top}(2,3),R^{\bot}(3,4)\}$ which is in the closed-world form.
The data complexity of WFOMC is extended to the case with evidence, where the evidence is treated as another input (besides the domain size) of the problem. Note that the number of ground literals in the evidence is polynomial in the domain size. Given a class $\mathfrak{E}$ of evidence, we say that a first-order fragment with $\mathfrak{E}$ is domain-liftable if the WFOMC conditioned on any evidence in $\mathfrak{E}$ can be computed in time polynomial in the domain size.
@@
The Gaifman graph with respect to the domain $\Delta$ and the binary evidence $\mathcal{E}$ is an undirected graph $G_{\Delta,\mathcal{E}}$ of $n$ vertices, in which there is an edge between $a$ and $b$ if and only if $\mathcal{E}$ contains $P(a,b)$ or $P(b,a)$ for some $P\in\mathcal{P}$.
Figure 1 shows an example of the Gaifman graph of the closed-world binary evidence
$\{R(a,b),R(a,c),R(b,c),R(c,d),R(c,e),R(d,g),R(e,g)\}$.
@@ 3.3 Implications of the Result
Our algorithm could be also applied to asymmetric WFOMC problems, whose details we defer to the full version due to space constraints.
@@ 4 Experiments
We evaluated the performance of our algorithm [^2] Our implementation can be found in: https://github.com/kulavacl/WFOMC---binary-evidence on several examples and compared it with two other model counting algorithms, d4333https://github.com/crillab/d4 and Forclift [^4] https://github.com/UCLA-StarAI/Forclift. d4 is a propositional model counter that compiles a formula in conjunctive normal form to deterministic decomposable negation normal form (d-DNNF). It generally performs well when compared to other model counters, but it cannot leverage lifted methods. Forclift is a model counter similarly to our algorithm that computes lifted inference and accepts binary evidence. All experiments were run on a laptop with Intel i7-10510U CPU and 32GB of RAM.
@@ 4.1 Friends and Smokers
In our experiments, we provide open-world evidence that corresponds to the persons coming in cliques of 3 (i.e., due to the open-world nature of the evidence, they may also have other friends beyond the cliques), and suppose $|\Delta|=3\cdot k$ ($k\geq 1$). We compared the runtime of our algorithm (labeled TD-WFOMC), d4 and Forclift. Figure 2 shows that our algorithm runs significantly faster than the other two, and scales well to large domain size.
@@ 4.2 Watts-Strogatz Model
Another interesting problem we can tackle using WFOMC is performing inference on Watts-Strogatz (WS) model (watts_collective_1998), which is a random graph model widely used in network science to generate small-world networks.
The generation process of a random graph starts with a regular ring lattice with $N$ nodes, where each node is connected to exactly $K$ other nodes (assuming $K$ is an even integer), $K/2$ on each side, warping over the end. To generate the random graph, each edge $(i,j)$ is rewired with probability $\beta$ in some order, meaning that edge $(i,j)$ is replaced with edge $(i,k)$ ($i\neq k$) such that edge $(i,k)$ does not exist already. The Watts-Strogatz model can be encoded as a Markov Logic Network (MLN) $\Psi$:
@@
$\infty,$ $\forall x\forall y:Edge(x,y)\leftrightarrow(WiredEdge(x,y)\vee WiredEdge(y,x)),$
@@ 5 New Results on the Stable Seating Arrangement Problem
Let $SSA(\mathcal{A},\mathcal{P},G,\mathcal{K})$ and $EFSA(\mathcal{A},\mathcal{P},G,\mathcal{K})$ be the counting version of the stable and envy-free seating arrangements problems respectively. We first show that there is an $\text{FO}^{2}$ sentence $\Psi$, unary evidence $\mathcal{U}$, binary evidence $\mathcal{E}$, and a set of cardinality constraints $\mathcal{C}$, such that $SSA(\mathcal{A},\mathcal{P},G,\mathcal{K})$ can be obtained from $\mathsf{WFOMC}(\Psi\wedge\mathcal{U}\wedge\mathcal{E}\wedge\mathcal{C},V_{G},w,\overline{w})$ in time polynomial in $|V_{G}|$.
@@
where $d_{a}$ is the degree of $a$. Recall that we label the outgoing edges of $a$ in such a way that $a$ is connected to each of its neighbors using a differently labeled edge.
The unary evidence $\mathcal{U}$ is only necessary for seating graphs with at least one vertex $a$ with degree $d_{a}<d$. In such case, we define

### v2

@@ 1 Introduction
We consider the task of computing WFOMC for $\text{FO}^{2}$ and $\text{C}^{2}$ with evidence, the interpretation of certain predicates. For example, consider the sentence
@@
over the domain $\Delta=\{1,2,3,4\}$. An evidence can be the specification that E(1,2),E(1,3),E(2,3),E(1,4)E(1,2),E(1,3),E(2,3),\\ E(1,4) and their reverse atoms are true and other atoms of $E$ are false. In this case, each model is an assignment of $I(1),\cdots,I(4)$ satisfying $\Psi$, which represents an independent set of the graph with vertices $\{1,2,3,4\}$ and undirected edges $(1,2),(1,3),(2,3),(1,4)$. The WFOMC of $\Psi$ with the above evidence over $\Delta$ with unit weight for each model equals to the number of independent sets on the graph.
@@ 1.1 Our Contributions
We propose an algorithm to compute WFOMC for an $\text{FO}^{2}$ or $\text{C}^{2}$ sentence with cardinality constraints, unary evidence and binary evidence whose Gaifman graph, the undirected graph with vertices corresponding to domain elements in which two vertices $a$ and $b$ are connected if and only if there is a binary predicate $P$ such that the truth value of $P(a,b)$ or $P(b,a)$ is specified in the binary evidence, is of bounded treewidth. The algorithm runs in time polynomial in the domain size. This offers a new approach to overcome the symmetry limitation in lifted inference, and expands the expressive power and applicability of WFOMC. The key idea of our algorithm is the dynamic programming for computing WFOMC for the universally quantified $\text{FO}^{2}$ sentence with evidence on the tree decomposition of the underlying Gaifman graph. Existing techniques help to extend the result to $\text{FO}^{2}$ and $\text{C}^{2}$ with cardinality constraints.
@@ 2.5 Evidence
The two forms of the evidence can be transformed to each other. Transforming the closed-world representation to the open-world representation is straightforward by adding the default negative ground literals to the evidence set. For instance, the closed-world evidence $\mathcal{E}=\{R(1,2),R(2,2)\}$ of the predicate set $\mathcal{P^{\prime}}=\{R\}$ over the domain $\Delta=\{1,2\}$ corresponds to the open-world evidence $\mathcal{E}^{\prime}=\{\lnot R(1,1),R(1,2),\lnot R(2,1),R(2,2)\}$. On the other hand, for example, consider the open-world evidence $\mathcal{E}=\{R(1,2),R(2,3),\lnot R(3,4)\}$. To transform $\mathcal{E}$ to the closed-world form, we introduce two fresh binary predicates $R^{\top}$ and $R^{\bot}$, add the sentence $\forall x\forall y:(R^{\top}(x,y)\to R(x,y))\land(R^{\bot}(x,y)\to\lnot R(x,y))$ and write $\mathcal{E}^{\prime}=\{R^{\top}(1,2),R^{\top}(2,3),R^{\bot}(3,4)\}$ which is in the closed-world form.
The data complexity of WFOMC is extended to the case with evidence, where the evidence is treated as another input (besides the domain size) of the problem. Note that the number of ground literals in the evidence is polynomial in the domain size. Given a class $\mathfrak{E}$ of evidence, we say that a first-order fragment $\mathfrak{L}$ with $\mathfrak{E}$ is domain-liftable if the WFOMC conditioned on any evidence in $\mathfrak{E}$, i.e., WFOMC of $\Psi\land\mathcal{E}$ for every $\Psi\in\mathfrak{L}$ and every $\mathcal{E}\in\mathfrak{E}$, can be computed in time polynomial in the domain size.
@@
The Gaifman graph with respect to the domain $\Delta$ and the closed-world binary evidence $\mathcal{E}$ is an undirected graph $G_{\Delta,\mathcal{E}}$ of $n$ vertices, in which there is an edge between $a$ and $b$ if and only if $\mathcal{E}$ contains $P(a,b)$ or $P(b,a)$ for some $P\in\mathcal{P}$.
Figure 1 shows an example of the Gaifman graph of the closed-world binary evidence $\{R(a,b),\\ R(a,c),R(b,c),R(c,d),R(c,e),R(d,g),R(e,g)\}$.
@@ 4 Experiments
We evaluated the performance of our algorithm on several examples and compared it with two other model counting algorithms, d4 (d4), Forclift (forclift), and Crane2 (crane). d4 is a propositional model counter that compiles a formula in conjunctive normal form to deterministic decomposable negation normal form (d-DNNF). It generally performs well when compared to other model counters, but it cannot leverage lifted methods. Forclift is a model counter similarly to our algorithm that computes lifted inference and accepts binary evidence. Crane2 is a model counter based on Forclift, which performs very well on a limited set of problems, but does not support full range of $\text{FO}^{2}$ sentences yet. For this reason, Crane2 is only used for friends and smokers experiments. All experiments were run on a laptop with Intel i7-10510U CPU and 32GB of RAM.
@@ 4.1 Friends and Smokers
In our experiments, we provide open-world evidence that corresponds to the persons coming in cliques of 3 (i.e., due to the open-world nature of the evidence, they may also have other friends beyond the cliques), and suppose $|\Delta|=3\cdot k$ ($k\geq 1$). We compared the runtime of our algorithm (labeled TD-WFOMC), d4, Forclift, and Crane2. Figure 2 shows that our algorithm runs significantly faster than the other three, and scales well to large domain size. We forgo the symmetry of friendships for Crane2 as symmetry is not supported. In Figure 3, we report the runtime of our algorithm for fixed domain size of 60 while varying the sizes of friend cliques.
@@
Figure 3: Time to compute the model count of friends and smokers problem with binary evidence, fixing $|\Delta|=60$ for varying size of friend cliques.
@@ 4.2 Watts-Strogatz Model
Another interesting problem we can tackle using WFOMC is performing inference on Watts-Strogatz (WS) model (watts_collective_1998), which is a random graph model widely used in network science to generate small-world networks. The generation process of a random graph starts with a regular ring lattice with $N$ nodes, where each node is connected to exactly $K$ other nodes (assuming $K$ is an even integer), $K/2$ on each side, warping over the end. To generate the random graph, each edge $(i,j)$ is rewired with probability $\beta$ in some order, meaning that edge $(i,j)$ is replaced with edge $(i,k)$ ($i\neq k$) such that edge $(i,k)$ does not exist already. The Watts-Strogatz model can be encoded as an MLN $\Psi$:
@@
$\infty,$ $\forall x\forall y:Edge(x,y)\leftrightarrow$
$(WiredEdge(x,y)\vee WiredEdge(y,x)),$
@@ 5 New Results on the Stable Seating Arrangement Problem
Let $SSA(\mathcal{A},\mathcal{P},G,\mathcal{K})$ and $EFSA(\mathcal{A},\mathcal{P},G,\mathcal{K})$ be the counting version of the stable and envy-free seating arrangements problems respectively. We first show that there is an $\text{FO}^{2}$ sentence $\Psi$, unary evidence $\mathcal{U}$, binary evidence $\mathcal{E}$, and a set of cardinality constraints $\mathcal{C}$, such that $SSA(\mathcal{A},\\ \mathcal{P},G,\mathcal{K})$ can be obtained from $\mathsf{WFOMC}(\Psi\wedge\mathcal{U}\wedge\mathcal{E}\wedge\mathcal{C},V_{G},w,\overline{w})$ in time polynomial in $|V_{G}|$.
@@
where $d_{a}$ is the degree of $a$. Recall that we label the outgoing edges of $a$ in such a way that $a$ is connected to each of its neighbors using a differently labeled edge. The unary evidence $\mathcal{U}$ is only necessary for seating graphs with at least one vertex $a$ with degree $d_{a}<d$. In such case, we define
@@ 6 Conclusion
Acknowledgements
Václav Kůla is supported by the Central Europe Leuven Strategic Alliance (CELSA) project Towards Scalable Algorithms for Neuro-Symbolic AI. Yuanhong Wang is supported by National Natural Science Foundation of China (No.62506141). Ondřej Kuželka is supported by the Czech Science Foundation project 23-07299S (Statistical Relational Learning in Dynamic Domains).